OpenStudy (anonymous):

Subtract. Write your answer in lowest terms.29/5 - 4/5 A) A) 5 B) B) 33/5 C) (C) 228/5 D) D) None of the other choices.

OpenStudy (anonymous):

\[\frac{ 29 }{ 5 } - \frac{ 4 }{ 5 } = \frac{ 29 - 4 }{ 5 } =\]